OpenStudy (chloee.in16):

Make a conjecture about the next item in the sequence. 1, 4, 16, 64, 256 Select one: a. 1024 b. 1025 c. 4096 d. 1022

OpenStudy (michele_laino):

hint: it is a geometrical sequence whose first term is 4 and the constant is 4, so the 6th term is: \[{a_6} = {a_1}{q^{6 - 1}} = 4 \cdot {4^5} = ...?\]

OpenStudy (michele_laino):

oops.. geometric*

OpenStudy (michele_laino):

fist term is 1

OpenStudy (michele_laino):

so we get: \[{a_6} = {a_1}{q^{6 - 1}} = 1 \cdot {4^5} = ...?\]

OpenStudy (michele_laino):

first* term

OpenStudy (er.mohd.amir):

2^0,2^2,2^4,2^6,2^8,2^10 so 2^10 is ans

OpenStudy (agent0smith):

All you need to do is figure out the pattern: 1, 4, 16, 64, 256 the pattern being that each number is multiplied by 4, to get the next number.

OpenStudy (chloee.in16):

Alright! thank you all so much! (:

OpenStudy (michele_laino):

I think that we can note this: \[\frac{{256}}{{64}} = \frac{{64}}{{16}} = \frac{{16}}{4} = \frac{4}{1} = 4\]

OpenStudy (er.mohd.amir):

its depends on how u see patter in Q but main think is that ans is correct at end